Can you please help me solve this problem?
"Nine plus twenty divided by an unknown number equals eleven"
I keep getting one-tenth. Thanks!

Answer by rfer(16322)

9+20/x=11
20/x=11-9
20/x=2
20=2x
20/2=x
10=x